OpenStudy (anonymous):

The integral of (tan^4xsec^2x)dx Help me!!!

OpenStudy (anonymous):

you can use u-substitution with the tan and sec^2x. the derivative of tanx=sec^2x

OpenStudy (anonymous):

Okay what should dv be then?

OpenStudy (anonymous):

dv=dx

OpenStudy (anonymous):

u=tan(x) du=sec^2(x)dx

OpenStudy (anonymous):

Okay so you are just gonna use U-substitution not anything else?

OpenStudy (anonymous):

yes, thats how i would approach the problem

OpenStudy (anonymous):

So the answer is (tan^2x)/2+c?

OpenStudy (anonymous):

But do you know how to figure it out and work it out?

OpenStudy (anonymous):

Yes, I do :) Thanks :P

OpenStudy (anonymous):

thats not correct, if you wolframalpha it you get (tan^5(x))/5 +c

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+%28tan%5E4xsec%5E2x%29dx This site is a great way to check your answers